80 g of hydrogen peroxide, H2 O2 (aq) react and give water and oxygen gas. 1. write the balanced reaction equation for the reaction. 2. how many grams of water is made?


1
Expert's answer
2021-12-08T15:23:01-0500

reaction equation

2H2O2"\implies" 2H2O + O2

moles of water formed?

ratio of reaction above is

2 : 2 : 1 respectively for peroxide, water and oxygen

moles of peroxide reacting = mass / molecular mass

= 80/ 2+32= 2.353

moles of water = moles of hydrogen peroxide

mass of water = moles * molar mass

= 2.353 * (2 + 16)

ans = 42.4 grams